OpenStudy (anonymous):

-5/2(w-1)=15 what's w?

OpenStudy (helder_edwin):

multiply by -2/5 and then add +1

OpenStudy (anonymous):

huh? could you show me please

OpenStudy (anonymous):

don't u muliply 15/2 by w and -1 first?

OpenStudy (anonymous):

i mean by -5/2

OpenStudy (helder_edwin):

\[ \large -\frac{5}{2}(w-1)=15 \] \[ \large w-1=15\cdot\frac{-2}{5}=-6 \] \[ \large w=-6+1=-5 \]

OpenStudy (helder_edwin):

yes u could do that. u should get the same solution

OpenStudy (anonymous):

thank u!! i like it ur way better (:

OpenStudy (helder_edwin):

u r welcome